# Defense Strategies for Autonomous Multi-agent Systems: Ensuring Safety and Resilience Under Exponentially Unbounded FDI Attacks

### v1

@@ III Safe and Resilient Controller Design
where $\zeta_{i}$ is the local state on the OL (the false data injected into the bus, on which the data driving the observer dynamics is transmitted) and $\gamma^{ol}_{i}$ represents the attack signal on the OL, $\vartheta_{i}$ is adaptively tuned by (11) with constant $q_{i}>0$, and $\xi_{i}$ represents the gathered neighborhood relative information on the OL given by
@@
$H_{i}=\Pi_{i}-K_{i}.$ (15)
The matrix $P_{i}$ is obtained by solving the following algebraic Riccati equation[6]:
@@
In order to guarantee the safety constraints to avoid collision, we then design the safety-aware control input $u^{s}_{i}$, which is computed by solving the following optimization problem:
$\min_{u^{s}_{i}}\|u^{s}_{i}-u_{i}^{c}\|^{2},$ (17)
subject to the constraints:
$\inf_{u_{i}\in\mathcal{U}_{i}}\left[L_{f}h_{S_{ij}}+L_{g}h_{S_{ij}}u_{i}\right]\leqslant-\delta_{ij}h_{S_{ij}}-2(x_{i}-x_{j})^{\top}\dot{x}_{j},$ (18)
where $\delta_{ij}$ is the parameters which influences the intensity of the safety constraints. We select $\alpha(h_{S_{ij}})=\delta_{ij}h_{S_{ij}}$[22]. Let $\Delta u_{i}\,\triangleq\,u^{s}_{i}-u_{i}^{c}$.
Remark 3.
In [23], the scalar coupling design relies on knowledge of the global graph topology. However, as shown in (11), the adaptive coupling gain $\vartheta_{i}$ is designed without requiring this global topology information. Therefore, the controller is fully distributed.
Remark 4.
The optimization problem formulated above ensures that the control input not only respects the safety constraints but also minimizes the deviation from the conventional control input. The solution is obtained by solving the QP problem [24]. The constraints imposed confine the optimizer’s solution $u_{i}^{s}$ within a specific region around $u_{i}^{c}$, thereby rendering $\Delta u_{i}$ bounded [11].
The SAAR control system is shown in Fig. 1. To address the attacks in (2), we finally design the following control input signal
$u_{i}$ $=u^{s}_{i}-\hat{\gamma}^{a}_{i},$ (19)
@@
where the compensational signal $\hat{\gamma}^{a}_{i}$ mitigates the adverse effects caused by the attack $\gamma^{a}_{i}$ (the false data injected into the communication channel, on which the data transmitted from the controller to the actuator is carried) and $\gamma^{ol}_{i}$, and $\hat{\rho}_{i}$ is adaptively tuned by (21) with constant $\alpha_{i}>0$.
Figure 1: Block diagram of the SAAR control system.
To facilitate the stability analysis, we define the follower-observer tracking error vector as
$\varepsilon\,\triangleq\,x-\zeta,$ (22)
where $\zeta={[{\zeta_{1}^{\top},...,\zeta_{N}^{\top}}]^{\top}}$. Define the observer containment error vector as
@@
Given Assumptions 1- 4, considering the heterogeneous autonomous MAS composed of (1)-(3) in the presence of EU-FDI attacks on both CIL and OL, the SAAR control problem is solved by designing the fully-distributed safe and resilient controller consisting of (10)-(21).
Proof: Note that the global form of (12) is
$\begin{gathered}\xi=-\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{N})\big{(}\zeta-\bar{x}_{r}\big{)}\hfill\\ =-\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{N})\Delta_{o},\hfill\end{gathered}$ (24)
where $\xi={[{\xi_{1}^{\top},...,\xi_{N}^{\top}}]^{\top}}$. Based on Lemma 3, since $\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{n})$ is nonsingular , to prove that $\Delta_{o}$ is UUB is equivalent to proving that $\xi$ is UUB. The vector form of $\dot{\zeta}_{i}$ in (10) is
@@
$\begin{gathered}\dot{\xi}=-\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{n})\big{(}\dot{\zeta}-\dot{\bar{x}}_{r}\big{)}\hfill\\ =-\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{n})\big{(}(I_{N}\otimes S)\zeta+\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{n}}\big{)}\xi\hfill\\ +\gamma^{ol}-(I_{N}\otimes S)\bar{x}_{r}\big{)}\hfill\\ =(I_{N}\otimes S)\xi-\sum_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{n})\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{n}}\big{)}\xi\hfill\\ -\sum_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{n})\gamma^{ol}.\hfill\end{gathered}$ (26)
@@
$V^{{}^{\prime}}=\frac{1}{2}\sum_{i=1}^{N}\xi_{i}^{\top}\xi_{i}\exp(\vartheta_{i}).$ (27)
The time derivative of $V^{{}^{\prime}}$ along the trajectory of (26) is given by
$\begin{gathered}\dot{V}^{{}^{\prime}}=\sum\limits_{i=1}^{N}\big{(}\xi_{i}^{\top}{\dot{\xi}_{i}}\exp(\vartheta_{i})+\frac{1}{2}\xi^{T}_{i}\xi_{i}\exp(\vartheta_{i})\dot{\vartheta}_{i}\big{)}\hfill\\ =\xi^{\top}\operatorname{diag}\big{(}\exp(\vartheta_{i})\otimes I_{N}\big{)}\dot{\xi}+\frac{1}{2}\xi^{T}_{i}\big{(}\operatorname{diag}(\exp(\vartheta_{i})\dot{\vartheta}_{i})\otimes I_{N}\big{)}\hfill\\ \times\xi\hfill\\ =\xi^{\top}\operatorname{diag}\big{(}\exp(\vartheta_{i})\otimes I_{N}\big{)}\bigg{(}(I_{N}\otimes S)\xi-\sum_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{N})\hfill\\ \times\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi-\sum_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{N})\gamma^{ol}\bigg{)}+\frac{1}{2}\xi^{\top}\hfill\\ \times\big{(}\operatorname{diag}(\dot{\vartheta}_{i})\otimes I_{N}\big{)}\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes I_{N}\big{)}\xi\hfill\\ \leqslant\sigma_{\max}(S)\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\|\xi\|-\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\hfill\\ \times\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|^{2}+\sigma_{\max}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\hfill\\ \times\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\|\gamma^{ol}\|+\frac{1}{2}\max_{i}(\dot{\vartheta}_{i})\hfill\\ \times\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\|\xi\|\hfill\\ =-\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\hfill\\ \times\bigg{(}\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|-\sigma_{\max}(S)\hfill\\ /\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\|\xi\|-\sigma_{\max}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\hfill\\ \times\|\gamma^{ol}\|-\frac{1}{2}\max_{i}(\dot{\vartheta}_{i})/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\|\xi\|\bigg{)}.\hfill\\ \end{gathered}$ (28)
For convenience, denote $\phi_{a}=\sigma_{\max}(S)/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}$ and $\phi_{b}=\sigma_{\max}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}$, which are both positive constants. To let $\dot{V}^{{}^{\prime}}\leqslant 0$, we need
@@
By LaSalle’s invariance principle [25], $\xi_{i}$ is UUB. Therefore, $\Delta_{o}$ is UUB.
Next, we prove that follower-observer tracking error $\varepsilon_{i}$ is UUB. From (1), (5), (10), (19) and (15), we obtain the time derivative of the elements in (22) as

### v2

@@ III Main Result of SAAR Controller Design
where $\zeta_{i}$ is the local state on the OL and $\gamma^{ol}_{i}$ represents the attack signal on the OL, $\vartheta_{i}$ is adaptively tuned by (10) with constant $q_{i}>0$, and $\xi_{i}$ represents the gathered neighborhood relative information on the OL given by
@@
$H_{i}=\Pi_{i}-K_{i},$ (14)
where the matrix $P_{i}$ is obtained by solving the following algebraic Riccati equation[7]:
@@
Define the follower-observer tracking error as
$\varepsilon_{i}=x_{i}-\zeta_{i}.$ (16)
We consider the additional EU-FDI attack signal in the CIL for each follower, $\gamma^{a}_{i}$. To address these attacks, we then design the following attack-resilient control input signal
$u^{r}_{i}$ $=u^{c}_{i}-\hat{\gamma}^{a}_{i},$ (17)
@@
where the compensational signal $\hat{\gamma}^{a}_{i}$ mitigates the adverse effects caused by the attack $\gamma^{a}_{i}$, and $\hat{\rho}_{i}$ is adaptively tuned by (19) with constant $\alpha_{i}>0$. Now the corrupted attack-resilient controller becomes $\bar{u}_{i}=u^{r}_{i}+\gamma^{a}_{i}$.
Define the observer containment error vector as
@@
Subsequently, we integrate the attack-resilience with collision-free safety as follows.
Lemma 3.
$\begin{gathered}\inf_{u_{i}\in\mathcal{U}_{i}}\left[L_{f_{i}}h_{S_{ij}}+L_{g_{i}}h_{S_{ij}}u_{i}\right]\leqslant-\alpha(h_{S_{ij}})\hfill\\ -2(x_{i}-x_{j})^{\top}A_{j}x_{j}-2(x_{i}-x_{j})^{\top}B_{j}u_{j},\end{gathered}$ (21)
Proof: By definition, the specific CBF is given as: $h_{S_{ij}}(x_{i},x_{j})=d_{s}^{2}-\|x_{i}-x_{j}\|^{2}.$ Taking the time derivative of $h_{S_{ij}}(x_{i},x_{j})$ along the system dynamics, we have: $\dot{h}_{S_{ij}}(x_{i},x_{j})=-\frac{d}{dt}\|x_{i}-x_{j}\|^{2}.$ Expanding $\|x_{i}-x_{j}\|^{2}=(x_{i}-x_{j})^{\top}(x_{i}-x_{j})$, its time derivative is $\frac{d}{dt}\|x_{i}-x_{j}\|^{2}=2(x_{i}-x_{j})^{\top}(\dot{x}_{i}-\dot{x}_{j}).$ Substituting this into $\dot{h}_{S_{ij}}(x_{i},x_{j})$, we obtain: $\dot{h}_{S_{ij}}(x_{i},x_{j})=-2(x_{i}-x_{j})^{\top}(\dot{x}_{i}-\dot{x}_{j}).$ From the system dynamics, we know: $\dot{x}_{i}=A_{i}x_{i}+B_{i}u_{i},\quad\dot{x}_{j}=A_{j}x_{j}+B_{j}u_{j}.$ Substituting $\dot{x}_{i}$ and $\dot{x}_{j}$ into the expression for $\dot{h}_{S_{ij}}(x_{i},x_{j})$, we get: $\dot{h}_{S_{ij}}(x_{i},x_{j})=-2(x_{i}-x_{j})^{\top}\left[(A_{i}x_{i}+B_{i}u_{i})-(A_{j}x_{j}+B_{j}u_{j})\right].$ Using the Lie derivative notation, this can be written as: $\dot{h}_{S_{ij}}(x_{i},x_{j})=L_{f_{i}}h_{S_{ij}}+L_{g_{i}}h_{S_{ij}}u_{i}+2(x_{i}-x_{j})^{\top}A_{j}x_{j}+2(x_{i}-x_{j})^{\top}B_{j}u_{j},$ where, $L_{f_{i}}h_{S_{ij}}=-2(x_{i}-x_{j})^{\top}A_{i}x_{i},L_{g_{i}}h_{S_{ij}}=-2(x_{i}-x_{j})^{\top}B_{i}$. Based on [21], to ensure forward invariance of $S_{S_{ij}}$, we impose the condition:
$\dot{h}_{S_{ij}}(x_{i},x_{j})\leqslant-\alpha(h_{S_{ij}}).$
Substituting the expanded form of $\dot{h}_{S_{ij}}(x_{i},x_{j})$, we obtain: $L_{f_{i}}h_{S_{ij}}+L_{g_{i}}h_{S_{ij}}u_{i}\leqslant-\alpha(h_{S_{ij}})-2(x_{i}-x_{j})^{\top}A_{j}x_{j}-2(x_{i}-x_{j})^{\top}B_{j}u_{j}.$ This inequality ensures that the safe set $S_{S_{ij}}$ is forward invariant. Thus, if the control input $u_{i}$ satisfies the optimization criterion:
$\inf_{u_{i}\in\mathcal{U}_{i}}\left[L_{f_{i}}h_{S_{ij}}+L_{g_{i}}h_{S_{ij}}u_{i}\right]\leqslant-\alpha(h_{S_{ij}})-2(x_{i}-x_{j})^{\top}A_{j}x_{j}-2(x_{i}-x_{j})^{\top}B_{j}u_{j},$ then $\|x_{i}-x_{j}\|\geqslant d_{s}$ is guaranteed for all $t\geqslant 0.\hfill\blacksquare$
Figure 1: Pairs.
For control affine systems, CBFs lead to linear constraints on the control inputs $u_{i}$ that can be enforced online. To ensure safety constraints and prevent collisions, we design the SAAR control input $u_{i}$ by solving the following QP. This optimization minimizes the difference between the actual SAAR control and the corrupted attack-resilient control input $\bar{u}_{i}$, while enforcing the constraints:
$\min_{u_{i}}\|u_{i}-\bar{u}_{i}\|^{2},$ (22)
subject to the constraints:
$\begin{gathered}\inf_{u_{i}\in\mathcal{U}_{i}}\left[L_{f_{i}}h_{S_{ij}}+L_{g_{i}}h_{S_{ij}}u_{i}\right]\leqslant-\delta_{ij}h_{S_{ij}}\hfill\\ -2(x_{i}-x_{j})^{\top}A_{j}x_{j}-2(x_{i}-x_{j})^{\top}B_{j}u_{j},\end{gathered}$ (23)
where $1\leqslant i<j<N$, which ensures the pairwise distances are defined for all agent pairs $(i,j)$ and each pair is counted exactly once as shown in Fig 1. $\delta_{ij}$ is the parameters which influences the intensity of the safety constraints. We select $\alpha(h_{S_{ij}})=\delta_{ij}h_{S_{ij}}$[22]. Let $\Delta u_{i}=u_{i}-\bar{u}_{i}$, which is obtained by solving a convex QP with linear constraints over the feasible set $\mathcal{U}_{i}$ (which is assumed to be compact or otherwise bounded in practice). Therefore, $\Delta u_{i}$ is bounded. The efficient solution of the QP problem enables the algorithm to be implemented in real-time [23]. Collision avoidance among agents is enforced through a sequential backward-constrained optimization scheme. Starting from the highest-indexed agent pair, the solution for each agent $i$ is computed using constrained optimization based on results from previously solved pairs $(i,j)$, where $j>i$. For example, if the total number of followers $N=4$, the resulting pairs are $(1,2)$, $(1,3)$, $(1,4)$, $(2,3)$, $(2,4)$, and $(3,4)$, fully enumerating all distinct pairwise interactions. In the updating algorithm, inputs for agents are sequentially updated by first addressing pairs involving the highest-indexed agents. Specifically, the input of the highest-indexed agent remains unmodified, i.e., $u_{4}=\bar{u}_{4}$, the input for agent 3 is updated based on the $(3,4)$ pair. Then, using the updated information for agents 3 and 4, the input for agent 2 is updated through pairs $(2,3)$ and $(2,4)$. Finally, inputs for agent 1 are updated simultaneously using the pairs $(1,2)$, $(1,3)$, and $(1,4)$, incorporating previously updated inputs. An alternative approach from [21] with reduced neighborhoods on a disk graph provides a scalable solution with less computational cost.
The proposed SAAR control system is shown in Fig. 2.
Figure 2: Block diagram of the SAAR control system.
@@
Given Assumptions 1- 4, considering the heterogeneous MAS composed of (1)-(2) in the presence of EU-FDI attacks on both CIL and OL, the SAAR control problem is solved by designing the fully-distributed safe and resilient controller consisting of (9)-(19) and (22)-(23).
Proof: The global form of (11) is
$\begin{gathered}\xi=-\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{N})\big{(}\zeta-\bar{x}_{r}\big{)}=-\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{N})\Delta_{o},\hfill\end{gathered}$ (24)
where $\xi={[{\xi_{1}^{\top},...,\xi_{N}^{\top}}]^{\top}}$. Based on Lemma 2, $\sum_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{n})$ is nonsingular. To prove $\Delta_{o}$ is UUB is equivalent to proving that $\xi$ is UUB. The global form of $\dot{\zeta}_{i}$ in (9) is
@@
$\begin{gathered}\dot{\xi}=-\sum\nolimits_{\nu\in\mathscr{L}}(\Phi_{\nu}\otimes I_{n})\big{(}(I_{N}\otimes S)\zeta+\hfill\\ \big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{n}}\big{)}\xi+\gamma^{ol}-(I_{N}\otimes S)\bar{x}_{r}\big{)}\hfill\\ =(I_{N}\otimes S)\xi-\sum_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{n})\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{n}}\big{)}\xi\hfill\\ -\sum\nolimits_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{n})\gamma^{ol}.\hfill\end{gathered}$ (26)
@@
$V^{{}^{\prime}}=\tfrac{1}{2}\sum\nolimits_{i\in\mathscr{F}}\xi_{i}^{\top}\xi_{i}\exp(\vartheta_{i}).$ (27)
The time derivative of $V^{{}^{\prime}}$ along the trajectory of (26) is
$\begin{gathered}\dot{V}^{{}^{\prime}}=\sum\nolimits_{i\in\mathscr{F}}\big{(}\xi_{i}^{\top}{\dot{\xi}_{i}}\exp(\vartheta_{i})+\frac{1}{2}\xi^{T}_{i}\xi_{i}\exp(\vartheta_{i})\dot{\vartheta}_{i}\big{)}\hfill\\ =\xi^{\top}\operatorname{diag}\big{(}\exp(\vartheta_{i})\otimes I_{N}\big{)}\dot{\xi}+\frac{1}{2}\xi^{T}_{i}\big{(}\operatorname{diag}(\exp(\vartheta_{i})\dot{\vartheta}_{i})\otimes I_{N}\big{)}\xi\hfill\\ =\xi^{\top}\operatorname{diag}\big{(}\exp(\vartheta_{i})\otimes I_{N}\big{)}\bigg{(}(I_{N}\otimes S)\xi-\sum_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{N})\hfill\\ \times\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi-\sum_{r\in\mathscr{L}}(\Phi_{r}\otimes I_{N})\gamma^{ol}\bigg{)}+\frac{1}{2}\xi^{\top}\hfill\\ \times\big{(}\operatorname{diag}(\dot{\vartheta}_{i})\otimes I_{N}\big{)}\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes I_{N}\big{)}\xi\hfill\\ \leqslant\sigma_{\max}(S)\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\|\xi\|-\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\hfill\\ \times\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|^{2}+\sigma_{\max}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\hfill\\ \times\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\|\gamma^{ol}\|+\frac{1}{2}\max_{i}(\dot{\vartheta}_{i})\hfill\\ \times\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\|\xi\|\hfill\\ =-\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|\hfill\\ \times\bigg{(}\|\big{(}\operatorname{diag}(\exp(\vartheta_{i}))\otimes{I_{N}}\big{)}\xi\|-\sigma_{\max}(S)\hfill\\ /\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\|\xi\|-\sigma_{\max}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\hfill\\ \times\|\gamma^{ol}\|-\frac{1}{2}\max_{i}(\dot{\vartheta}_{i})/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}\|\xi\|\bigg{)}.\hfill\\ \end{gathered}$
Denote $\phi_{a}=\sigma_{\max}(S)/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}$ and $\phi_{b}=\sigma_{\max}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}/\sigma_{\min}\big{(}\sum_{r\in\mathscr{L}}\Phi_{r}\big{)}$, which are both positive constants. For $\dot{V}^{{}^{\prime}}\leqslant 0$, we need
@@
By LaSalle’s invariance principle [24], $\xi_{i}$ is UUB. Therefore, $\Delta_{o}$ is UUB. Next, we prove that follower-observer tracking error $\varepsilon_{i}$ is UUB. From (1), (4), (9), (17) and (14), we obtain the time derivative of (16) as